Magic Realm, Inc., has developed a new fantasy board game. The company sold 41,000 games last year at a selling price of $68 per game. Fixed expenses associated with the game total $738,000 per year, and variable expenses are $48 per game. Production of the game is entrusted to a printing contractor. Variable expenses consist mostly of payments to this contractor.

Required:

1-a. Prepare a contribution format income statement for the game last year.

1-b. Compute the degree of operating leverage.

2. Management is confident that the company can sell 49,610 games next year (an increase of 8,610 games, or 21%, over last year). Given this assumption:

a. What is the expected percentage increase in net operating income for next year?

b. What is the expected amount of net operating income for next year? (Do not prepare an income statement; use the degree of operating leverage to compute your answer.)
Answer #1

1a) Contribution margin income statement

Total Per unit
Sales 2788000 68
Variable cost 1968000 48
Contribution margin 820000 20
Fixed cost 738000
Operating income 82000

1b) Degree operating leverage = 820000/82000 = 10

2a) Net operating income increase = 21*10 = 210%

2b) Net operating income = 82000*3.1 = 254200